OpenStudy (anonymous):

verify identity of 1+cos2a = 2/ 1+tan^2a

OpenStudy (anonymous):

by (1) cos(2a) = cos²(a) - sin²(a) , (2) sin²(a) + cos²(a) = 1 , (3) tan²(a) + 1 = sec²(a) . Using (1) and (2) we get 1 + cos(2a) = 1 + cos²(a) - sin²(a) = 2 cos²(a) . cos²(a) = 1 / sec²(a) and applying (3), we find 2 cos²(a) = 2 / sec²(a) = 2 / (1 + tan²(a)